The relative denisty of a body is 2.5. What would be its mass if the same volume of water weighs 250g?

Grade:7

2 Answers

Pritam
6 years ago
Relative density=substance density÷ standard density(water)Hence R.d=2.5=M÷V But given that volume of water is same as that of solid &mass of water is given so equate it So volume will be cancel out &we getM=2.5×250. 1=m÷v (water), 1=250÷vM=625
Nitesh
6 years ago
Relative density = density of body ÷ density of water at 4°C2.5 = mass of body / mass of water (given volume of water = volume of body)2.5 × mass of water = mass of body =》 Mass of body = 2.5 × 250 gram =》 mass of body =625 gram